### v1

# Transition from positive to negative photoconductivity in AlGaN/GaN quantum-well heterostructures

Abstract
The AlGaN/GaN quantum-well heterostructures typically exhibit a positive photoconductivity (PPC) during the light illumination. Surprisingly, we found that introducing the GaN/AlN superlattice (SL) back barrier into N-polar AlGaN/GaN quantum-well heterostructures induces a transition in these heterostructures from PPC to negativie photoconductivity (NPC) as the SL period number increased at room temperature. This transition occurred under an infrared light illumination and can be well explained in terms of the excitation of hot electrons from the two-dimensional electron gas and subsequent trapping them in a SL structure. The NPC effect observed in N-polar AlGaN/GaN heterostructures with SL back barrier exhibits photoconductivity yield exceeding 85 and thus is the largest ones reported so far for semiconductors. In addition, NPC signal remains relatively stable at high temperatures up to 400 K. The obtained results can be interesting for the development of NPC related devices such as photoelectric logic gates, photoelectronic memory and infrared photodetectors.
I Introduction
The photoconductivity is a phenomenon in which the electric conductivity of a material increases upon the light illumination. In the literature, this effect is often called the positive photoconductivity (PPC) and is commonly observed in various semiconductors. However, in rare cases it is possible to observe the opposite effect to PPC, i.e. negative photoconductivity (NPC), in which the electric conductivity decreases upon illumination [1, 2, 3, 4, 5]. This anomalous effect was found in certain low-dimension materials, such as, nanowires,[6, 7, 8] MoS2 monolayers [9], quantum dots [10], nanorods [11], degenerated InN thin films [12], graphene [13], Van der Waals heterostructures [14, 15], as well as in a few bulk materials [16, 17, 18, 19, 20, 21] and InAs/InGaAs quantum well heterostructures [22]. The NPC effect found great potential applications in optoelectronic memory, low-power photodetectors, gas sensors etc. [23, 24, 25, 26, 27, 28, 29, 30] Furthermore, the combination of NPC and PPC effects can be used to build a photoelectric logic gate [31].
For almost 30 years, the AlGaN/GaN quantum well heterostructures were widely investigated because of their potential applications in fabricating various electrical and optical electronic devices such as field-effect transistors and photodetectors [32]. However, despite this fact experimental observations of the NPC effect in these heterostructures were very rare and limited mainly to a very low temperature [33], which hinders practical development of NPC devices based on the AlGaN/GaN heterostructures. In this work, we surprisingly discovered that introducing the GaN/AlN superlattice (SL) back barrier into N-polar AlGaN/GaN heterostructures caused the transition in these heterostructures from PPC to NPC effect as the SL period number increased at the room temperature during infrared light illumination. The observed NPC effect in the N-polar AlGaN/GaN quantum-well heterostructure with SL back barrier is the largest one reported for semiconductors so far - it exhibits the photoconductivity yield exceeding 85. We proposed the explanation of this phenomenon in terms of the excitation of hot electrons from the two-dimensional electron gas (2DEG) and subsequent trapping them in the SL structure.
II Sample structure and experiment
In this paper, we investigated the photoconductivity of the N-polar AlGaN/GaN heterostructures with GaN/AlN SL back barrier whose schematic illustration is shown in Fig. 1(a). The photoconductivity was analyses as a function GaN/AlN SL periods n, which was varying from 0 to 8 (see inset in Fig. 1(a)). The investigated structures were grown by metal-organic chemical vapor deposition (MOCVD) using trimethylgallium (TMGa), trimethylaluminum (TMAl), and ammonia as precursors on a (0001) sapphire substrate. The process stared from the growth of 100 nm AlN layer at 1025 oC followed by growth of 1 um unintentionally doped GaN buffer layer. Next, on the top of the GaN buffer layer a thin SL structure was deposited containing the alternating 10 nm thick GaN and 5 nm thick AlN layers as shown in the inset of Fig. 1(a)). Subsequently, a 30-nm-thick Al0.3Ga0.7N barrier layer was deposited on the SL structure followed by the growth of 20-nm thick undoped GaN channel layer. Finally, 3-nm-thick Al0.4Ga0.6N and 2-nm-thick GaN cap layers were deposited on the top of the structure. The details of the fabrication process and transmission electron microscopy (TEM) images of superlattice structures can be found in Ref. [34]. In the same Ref. 34, 2DEG density and electron mobility of the fabricated heterostructures, as obtained from the Hall transport measurements, can be also found. For the photocurrent measurements, the aluminum-titanium ohmic contacts were fabricated, as shown in Fig. 1(a). The current-voltage (I-V) characteristics proved good ohmic properties of contacts in all samples with different SL period numbers (see Fig. 1 (b)). The photoconductivity experiments were performed using a 150 W halogen lamp with filters of 430 nm, 610 nm and 990 nm wavelengths at room temperature and 400 K. The light intensity was kept at a relatively low level of 40 mW/cm2 to avoid sample heating. The illuminated area is indicated by the dashed red lines in the inset of Fig. 1 (b).
III Results and Discussion
III.1 Photocurrent data
Fig. 2 presents the transient photocurrent () measurements (light on/off) performed for N-polar AlGaN/GaN heterostructures with SL period =0, 2, 4, 6 and 8 at room-temperature under a bias voltage of 2 V and illumination with wavelengths: (a) 430 nm, (b) 610 nm and (c) 990 nm. The photocurrent, , is defined as the difference between the current under illumination () and in the dark () as schematically shown in the inset of Fig. 2(a).
In the case of the visible light illumination (Figs. 2 (a) and (b)) all samples exhibited the PPC effect. However, in the case of the infrared illumination (Fig. 2 (c)) a clear transition from PPC to NPC was observed with SL period increasing. In particular, the PPC effect occurred for the sample with =0 whereas for the samples with , the current decreased after the infrared light turn-on that indicated the NPC effect. It is interesting to note that for the sample with =2 there was no response at the infrared light illumination. In addition, for the sample with =4 the current suddenly decreased but for the samples with =6 and 8 this decreasing was less rapid. The strongest NPC effect, i.e. the largest current drop due to the illumination was observed for =4. For all samples where the NPC effect occurred, was far from the saturation within 60 s. However, in PPC, was saturated within 60 s for almost all cases. Furthermore, for the samples exhibiting NPC, the long persistent photoconductivity was observed, i.e. the current needed a long time to return to its initial dark value after the light turn-off. This long persistent NPC effect can be interesting for applications in the NPC optical memory devices. Contrary to NPC, the PPC decay is much faster after switching off the light (see Figs. 2(a)-(b) and (c) for samples with =0). This observation is consistent with the previous reports in which PPC decayed much faster than NPC [21].
In Figs. 3(a)-(c), we summarized the dependencies of the photocurrent from Fig. 2 as a function of the SL period number . In the case of 430 nm (Fig. 3(a)) weakly depends on up to =4. Above =4, decreased probably due to increasing the non-radiative state density [34]. The similar trend was also observed for a wavelength of 610 nm (Fig. 3(b)), except for =4 for which an enhancement of was registered. The reason of this increasing is unclear at present. Under the infrared light illumination (Fig. 3(c)) changed significantly for the samples between =0 and =4 with a clear change of sign at =2 from the positive to negative value. Above = 4, it seems that become saturated with . In addition it seems that there is no correlation bettwen the magnitude of the NPC effect and 2DEG density see Ref [34]. For the potential applications in the infrared photodetectors, we also verified the temperature impact on the observed NPC effect, as shown in Fig. 3(d). As can be seen from this figure, the temperature rising up to 400 K did not cause a significant decrease of NPC effect compared to the room temperature (see Fig. 2 (c) curve for =4). We can even say that the obtained NPC effect was relatively stable with temperature which is advantageous for applications in NPC devices. For the estimation of magnitude of the measured NPC, we calculated the photoconductivity yield defined as [21]:
where is the saturation current (after illumination). Because in our case NPC was not saturated within 60 s (see Fig. 2(c)), we extended the photocurrent transient measurements to much longer times, as show in Fig. 3(e). It is evident that starts to saturate around 5000 s but the full saturation is not reached within 40000 s. Nevertheless, this figure clearly indicates that the registered NPC effect is huge. In particular, at 40000 s, 0.85 mA, which is 85 of the dark current = 1 mA (for sample with =8, see Fig. 1(b) at bias 2 V). Therefore, we can assume that is 85 (see Eq. 1). For comparison, in the case of graphene of only 34 was reported. Note that the authors registered the full saturation of the photo-current after 7500 s. After this time the authors observed the current decreasing by 34 [13]. The highest was observed so far for Bi-doped MAPbBr3 perovskites. In this case, the photo-current was almost saturated after 1000 s illumination and of 67 was obtained [21]. Our significantly exceeds this value, which means that it is the highest ever reported for semiconductors. In this context, it should be highlighted that the large NPC effect with the value of 50-90 was also observed in the Van der Waals heterostructures [14, 15]. However, in those structures the high value 90 was achieved due to using the floating gate structure but not as an intrinsic material property [14]. Thus, we believe that our NPC effect is the highest one obtained for semiconductor materials.
III.2 Discussion
The origin of NPC phenomenon was attributed previously to generation of hot electrons, formation of metastable states (in particular the light induced trap states and DX centers), intraband transitions , surface plasmon scattering, trapping by deep-levels, adsorption of moisture and trion formation [9, 10, 21, 22, 35, 36]. From all of these hypotheses, the hot electron-based ones seem to be most suitable for understanding the occurrence of NPC effect in our AlGaN/GaN samples (just like it was observed in 2DEG InAs/InGaAs systems [22]). Below we will show that the observed NPC effect can be well explained in terms of hot electron generation and thermalization.
In our recent studies, we found from the Secondary Ion Mass Spectroscopy that the fabricated GaN/AlN SL may not contain the pure AlN layer but rather high Al-content AlGaN layer. On the other hand, such layer exhibits much lower thermal conductivity than pure AlN or GaN because of the dominant phonon-alloy scattering [37]. Thus, introducing the SL back barrier structure to the investigated N-polar AlGaN/GaN heterostructure should led to reducing the thermal conductivity. In other words, this means that the samples with the SL structure can be heated more effectively than the samples without SL. Next, upon the infrared radiation the hot electrons are excited from the quantum well at Al0.3Ga0.7N/GaN interface, as schematically shown in Fig. 4(a) (process 1). It should be noted that our recent band diagram simulation showed only one quantum well in the investigated N-polar AlGaN/GaN heterostructure (located at the Al0.3Ga0.7/GaN interface) when SL contains high Al-content AlGaN instead of pure AlN (see Fig. 4 in Ref. [34]). The excited hot electrons from the quantum well are subsequently trapped in the SL structure (see Fig. 4(a) process 3A) which causes that the energy of hot electrons is transferred to the host lattice leading to heating the near Al0.3Ga0.7/GaN interface due to the low thermal conductivity of the SL structure and thereby reducing the electron mobility in 2DEG. Beside the excitation electrons from the quantum well at Al0.3Ga0.7N/GaN interface, the infrared light must also lead to the excitation of electrons from deep-levels in the GaN buffer layer, as shown in Fig. 4 (process 4), since a relatively high PPC effect was observed in the structure without SL (see Fig. 2).
Based on the above processes, we can explain the observed NPC effect using the following approach. Firstly, we defined the dark current density, in our heterostructures. The main conductivity channel is 2DEG located at the Al0.3Ga0.7N/GaN interface. However, some contribution to the conductivity may also comes from the bulk electrons in the GaN buffer layer. Thus, is expressed as follows:
where is the 2DEG concentration at the Al0.3Ga0.7N/GaN interface, is the 2DEG mobility in the dark, is the concentration of electrons in the bulk and is the bulk electron mobility.
As we explained before, under the infrared illumination the electrons from the deep levels in the GaN buffer layer are excited to the conduction band and 2DEG mobility is reduced due to the hot electron excitation (see process 1 in Fig. 4 (a)). These processes modified the current as follows:
where is the 2DEG mobility in the case of light illumination and is the concentration of excess electrons excited from the deep levels in GaN buffer layer. Note that in Eq. 3 we reasonably assumed that the number of excited hot electrons is negligible compared to .From the combination of Eq. 2 and 3, we obtain the expression for the photocurrent as:
The trapped hot electrons in the SL structure (process 3A in Fig. 4(a)) lead to heating the near Al0.3Ga0.7N/GaN interface due to low thermal conductivity of SL and reducing 2DEG mobility. Then, if this reduction is sufficiently high, the following term in Eq. 4 will be larger than leading to the negative photocurrent ( 0) and thus the NPC effect. In the case of the sample without SL, the excited hot electrons from the quantum well are relaxed in the GaN buffer layer, as schematically shown in Fig. 4(b) (process 3B). This process does not lead to heating the near Al0.3Ga0.7N/GaN interface since the thick GaN buffer layer exhibits high thermal conductivity. Therefore, in the sample without SL and thus according to Eq. 4, we observed in this sample the PPC effect equal to .
The last issue is to clarify why the NPC effect was not observed under visible light illumination (Figs. 2(a) and (b)). This can be explained by the two hypothesis (here called A and B). The first hypothesis A based on the Stark-Einstein rule according to which transitions of electrons from the ground states to excited ones occur only if the energy of the incident photons match exactly the energy difference between these states. In our case, hot electrons appear due to electron excitations from the energy levels in 2DEG to unfilled bands in the conduction band as shown schematically in Fig. 5(a). The visible photons with specific wavelengths of 430 nm and 610 nm do not much well to such energy differences and may not lead to hot electron generation (see Fig. 5(b)). Instead of that all photons with the wavelengths of 430 nm and 610 nm are mostly absorbed at the deep-levels in the GaN buffer layer. This explains well why under visible light illumination a relatively strong PPC signal was observed (Figs. 2 (a) and (b)).
According to the hypothesis B, both visible and infrared light leads to the generation of hot-electrons from 2DEG. However, the optical cross section for excitations of electrons from the deep-level centers corresponding to the visible light spectrum is much higher than for excitations of electrons from 2DEG region. In consequence, most of the photons with wavelengths of 430 nm and 610 nm are absorbed by the deep-levels in the GaN layer and only few lead to the generation of hot-electrons from 2DEG, which causes that the PPC effect dominates (Figs. 2(a) and (b)). In contrast, for the deep-level centers corresponding to the infrared light spectrum the optical cross section can be relatively low which causes that most of infrared photons are absorbed in the 2DEG region leading to hot-electron excitation and the NPC effect (Fig. 2(c)).
III.3 Conclusions
We have experimentally observed the negative photoconductivity effect in the N-polar AlGaN/GaN quantum-well heterostructures after introduction a GaN/AlN superlattice back barrier. The NPC phenomenon appeared in the N-polar AlGaN/GaN quantum-well heterostructures with the SL back barrier under infrared light illumination at room-temperature. The obtained NPC exhibits the photoconductivity yield exceeding 85 which is the largest value reported so far for semiconductor materials. In addition, it remains relatively stable at high temperatures up to 400 K. We show that the observed NPC can be well explained in terms of the excitation of hot electrons from the 2DEG quantum well and subsequent trapping them in the SL structure. Our finding may offer a novel approach to design NPC devices based on the AlGaN/GaN quantum well heterostructures.
Acknowledgements.
The authors expresses gratitude to Norikazu Nakamura for his kind support and discussions The authors would like to thank Yoshiharu Kinoue and Takaaki Sakuyama for their support in experiments.References
- [1] R. Newman, Phys. Rev. 94, 278 (1954).
- [2] J. R. Haynes and W. Shockley, Phys. Rev. 81, 835 (1951).
- [3] W. C. Dunlap, Jr., Phys. Rev. 91, 1282 (1953).
- [4] C. M. Penchina, J. S. Moore, and N. Holonyak, Phys. Rev. 143, 634 (1966).
- [5] A. Berkeliev and K. Durdyev, Sov. Phys. Semicond. 5, 646 (1971).
- [6] H.M. Huang, R.S. Chen, H.Y. Chen, T.W. Liu, C.C. Kuo, C.P. Chen, H.C. Hsu, L.C. Chen, K.H. Chen, Y.J. Yang Appl. Phys. Lett., 96 (6) (2010).
- [7] L. Li, E. Auer, M. Liao, X. Fang, T. Zhai, U.K. Gautam, A. Lugstein, Y. Koide, Y. Bando, D. Golberg Nanoscale, 3 (3) (2011), pp. 1120-1126.
- [8] Y. Yang, X. Peng, H.S. Kim, T. Kim, S. Jeon, H.K. Kang, W. Choi, J. Song, Y.J. Doh, D. Yu Nano Lett., 15 (9) (2015), pp. 5875-5882.
- [9] Lui C H, Frenzel A J, Pilon D V, Lee Y H, Ling X, Akselrod G M, Kong J and Gedik N 2014 Phys. Rev. Lett. 113 166801.
- [10] A. Vardi, G. Bahir, S. E. Schacham, P. K. Kandaswamy and E. Monroy, J. Appl. Phys., 2010, 108.
- [11] Shuchi Kaushik, AshokKumar Kapoor, RohitKumar Pant, SaluruBaba Krupanidhi, Rajendra Singh Opt. Mater., 121 (2021), Article 111553
- [12] C. Wei, S. Chattopadhyay, M. De Yang, S.C. Tong, J.L. Shen, C.Y. Lu, H.C. Shih, L.C. Chen, K.H. Chen Phys. Rev. B Condens. Matter, 81 (4) (2010), pp. 1-5.
- [13] Biswas C, Gunes F, Duong D L, Lim S C, Jeong M S, Pribat D and Lee Y H 2011 Nano Lett. 11 4682.
- [14] Wang, Y. et al. ACS Nano 12, 9513–9520 (2018).
- [15] W. He, D. Wu, L. Kong, P. Yu, G. Yang, Adv. Mater. 2024, 36, 2312541.
- [16] C.M. Penchina, J.S. Moore, N. Holonyak Phys. Rev., 143 (2) (1966), pp. 634-636.
- [17] N. Kavasoglu, A.S. Kavasoglu, S. Oktik J. Phys. Chem. Solid., 70 (3–4) (2009), pp. 521-526.
- [18] Koide, Y.; Alvarez, J.; Imura, M.; Kleider, J.; Liao, M. Phys. Rev. B 2008, 78, 45112.
- [19] B.A. Akimov, V.A. Bogoyavlenskiy, L.I. Ryabova Phys. Rev. B Condens. Matter, 61 (23) (2000), pp 16045-16051
- [20] Balslev I 1966 Phys. Rev. 143 636
- [21] M. A. Haque, J.-L. Li, A. L. Abdelhady, M. I. Saidaminov, D. Baran, O. M. Bakr, S.-H. Wei, T. Wu, Adv. Optical Mater. 2019, 7, 1900865.
- [22] T. Akazaki, M. Yamaguchi, K. Tsumura, S. Nomura, H. Takayanagi Phys. E Low-dimens. Syst. Nanostruct., 40 (2008), pp. 1341-1343
- [23] Kim G, Kim I G, Baek J H and Kwon O K 2003 Appl. Phys. Lett. 83 1249.
- [24] Alvarenga D, Parra-Murillo C A, Penello G M, Kawabata R, Rodrigues W N, Miquita D R, Schmidt W, Guimarães P, Unterrainer K and Souza P L 2013 J. Appl. Phys. 113 043721
- [25] S S, Pires M P, Unterrainer K and Souza P L 2013 J. Appl. Phys. 113 043721.
- [26] Chaves A S and Chacham H 1995 Appl. Phys. Lett. 66 727.
- [27] Yang J, Li R, Huo N, Ma W L, Lu F, Fan C, Yang S, Wei Z, Li J and Li S S 2014 RSC Adv. 4 49873.
- [28] Liu W, Lee J S and Talapin D V 2013 J. Am. Chem. Soc. 135 1349.
- [29] Rahman S, Samanta S, Kuzmin A, Errandonea D, Saqib H, Brewe D L, Kim J, Lu J and Wang L 2019 Adv. Sci. 6 1901132.
- [30] Hassan M Y and Ang D S 2019 ACS Appl. Mater. Interfaces 11 42339.
- [31] Fan Z, Dutta D, Chien C J, Chen H Y, Brown E C, Chang P C and Lu J G 2006 Appl. Phys. Lett. 89 213110.
- [32] H. Morkoc, Handbook on Nitride Semicnductors and Devices, Vol. 3: GaN-based optical and eltronic devices (Wiley, Weinheim, 2009).
- [33] N. Biyikli; U. Ozgur; X. Ni; Y. Fu; H. Morkoc; Ç. Kurdak J. Appl. Phys. 100, 103702 (2006).
- [34] Matys, M., Yamada, A. and Ohki, T., Phys. Status Solidi RRL 2400379 (2025).
- [35] A. L. Powell, C. C. Button, J. S. Roberts, P. I. Rockett, H. G. Grimmeiss, and H. Pettersson, Phys. Rev. Lett. 67, 3010 (1991).
- [36] Naveen Kumar Tailor, Clara A. Aranda, Michael Saliba, and Soumitra Satapathi ACS Materials Letters 2022 4 (11), 2298-2320.
- [37] Weili Liu, Alexander A. Balandin, J. Appl. Phys. 97, 073710 (2005)